OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (campbell_st):

rewrite the equation and its \[7\sin^2(x) -14x +7 = 0\] divide each term by 7 \[\sin^2(x) -2\sin(x) + 1 = 0\] factor the perfect square then solve each factor for the solution.

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

it some times is easy to think of sin(x) is some variable a then 7a^2-14a+2=-5

OpenStudy (anonymous):

But what are all the possible solutions ?

OpenStudy (zzr0ck3r):

7a^2-14a+7=0 7a^2-7a-7a+7=0 7a(a-1)-7(a-1)=0 (7a-7)(a-1)=0 so a=1 s0 sin(x) = 1 when does sin(x) = 1?

OpenStudy (zzr0ck3r):

pi/2 +- 2pi*k where k runs through the integers.

OpenStudy (anonymous):

I am confused should that be \[7\sin^2x -14sinx +2=5\]?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

Yeah thats it

OpenStudy (anonymous):

If so it would be easiest to let \[u = sinx\] then the equation is\[7u^2 - 14u +2 = -5\]. Now add 5 to both sides of the equation and factor out the 7. This results in \[7u^2-14u+2=5\] \[7u^2-14u+7=0\] \[7(u^2-2u+1)=0\] \[u^2-2u+1=0\] This is the common quadratic trinomial that can be factored to \[(u-1)^2=0\] So we need to determine when u = 1. Since \[u = \sin x\], we are looking for any time where sin x = 1. This is when \[x = \frac{ \pi }{ 2 }+2\pi n\] where \[n \in \mathbb{Z}\]

OpenStudy (anonymous):

Substituting a variable in for a trig function to make it easier to solve and then re-substituting the trig function later after solving the equation is a common method of solving trigonometric equations. It is also used quite often in calculus.

OpenStudy (anonymous):

I hope that helps.

OpenStudy (anonymous):

Yeah thanks